# Diffusion models for multivariate subsurface generation and efficient probabilistic inversion

### v1

@@ 2.2 Diffusion Posterior Sampling (DPS)
In Bayesian inversion, one is interested in characterizing or sampling from the posterior pdf $p(\textbf{x}_{0}|\textbf{d})$ with the data vector d given by
@@
Unconditional modeling from $p(\textbf{x}_{0})$ can be achieved with a trained DM, using Eqs. 2 or 3. Following Bayes’ rule, sampling from the posterior pdf can be achieved using the score function (Song et al., 2021b)
@@
where the first term can be defined the trained neural network $S_{\theta}(\textbf{x}_{t},t)$ and $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ is a noise-dependent likelihood score. In the context of inverse modeling, we refer to $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$ as the prior score. If $S_{\theta}(\textbf{x}_{t},t)\simeq\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$ (Eq. 6), then $\hat{\textbf{x}}_{0}^{(t)}$ is derived using Tweedie’s formula (Eq. 5). Otherwise, it is directly obtained as $\hat{\textbf{x}}_{0}^{(t)}=S_{\theta}(\textbf{x}_{t},t)$ (Eq. 7). The likelihood score is given by an intractable integral
[…]
@@
In practical implementations in DM, the score in Eq. 12 is computed by backpropagating the log-likelihood values evaluated at each time step $t$.
This approach is referred to as Diffusion Posterior Sampling (DPS), and was proposed by Chung et al. (2023) as a general inference framework for both linear and nonlinear inverse problems. Assuming uncorrelated and homoscedastic Gaussian data errors, Chung et al. (2023) use the likelihood
$p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})=\frac{1}{\sqrt{(2\pi)^{n}\sigma_{\textbf{d}}^{n}}}\exp{\left(-\frac{||\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d}||_{2}^{2}}{2\sigma_{\textbf{d}}^{2}}\right)},$ (13)
@@
$\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\textbf{x}_{t})\simeq\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})=-\frac{1}{\sigma_{\textbf{d}}^{2}}\nabla_{\textbf{x}_{t}}||\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d}||_{2}^{2}.$ (14)
This approximation does not account for errors in the approximated $\hat{\textbf{x}}_{0}^{(t)}$. In their implementation, Chung et al. (2023) propose a weight to the likelihood score, replacing $\frac{1}{\sigma_{\textbf{d}}^{2}}$ with an hyperparameter $\rho$, to account for the impact of the likelihood approximation used. In their accompanying code, the authors further replace the sum of square error in Eq. 14 with the root-mean-square error (RMSE), justifying this choice as a solution for increased stability (see open review in Chung et al., 2023). Indeed, this arbitrary choice will drastically decrease the magnitude of the likelihood score and, hence, decrease the weight given to the data in the sampling process.
The conditional score approximation used within DPS results in a too low variance of the generated samples, and the sampling process has been rather associated to the maximization of a posterior (Xu et al., 2025). Improvements to this approximation should include the denoiser’s ($\hat{x}_{0}^{(t)}$) uncertainty (Song et al., 2023; Hamidi and Yang, 2025; Yismaw et al., 2025; Boys et al., 2024; Peng et al., 2024). Following Efron (2011), the denoiser’s variance can be computed using Tweedie’s formula for the second moment
@@
where $\nabla_{\textbf{x}_{t}}^{2}\log p_{t}(\textbf{x}_{t})$ is the Hessian matrix of the score of the data distribution $\textbf{x}_{t}$. Using the Hessian directly is computationally demanding, as its calculation scales exponentially with the model’s dimensionality (Boys et al., 2024). The Covariance-Aware Diffusion Posterior Sampling (Hamidi and Yang, 2025) for linear inversion uses an approximation of the Hessian, computed as the difference between the scores at consecutive and sufficiently small time-steps (i.e., using the finite difference method). The $\Pi$GDM (Song et al., 2023) and CoDPS (Yismaw et al., 2025) methods use time-step dependent heuristic approximations of the denoiser’s standard deviation for linear inverse problems, representing significantly faster solutions. Peng et al. (2024) derive a formulation for optimal covariance quantification for specific DM networks able to predict the conditional variance used in DDIM ($\tilde{\sigma_{t}}$ in Eq. 8). For more general applications, they suggest to estimate the denoiser variance directly from the denoising errors of the trained network, using known test data. In the following, we follow the latter approach, and introduce a general and consistent formulation for approximate posterior sampling for both linear and nonlinear conditioning.
@@ 2.2.1 Corrected Diffusion Posterior Sampling (CDPS)
The proposed corrected DPS (CDPS), in which the likelihood formulation accounts for the denoising-step dependent error $\epsilon_{\hat{\textbf{x}}_{0}}^{(t)}$ of the trained model. The $\epsilon_{\hat{\textbf{x}}_{0}}^{(t)}$ is the largest at early steps of the backward process and decrease with decreasing $\sigma_{\textbf{t}}$. The forward problem at time t can be written as (c.f., Eq. 9).
@@
Assuming an ideal denoiser, we assume $\epsilon_{\hat{\textbf{x}}_{0}}^{(t)}$ to be spatially uncorrelated and Gaussian-distributed (Song et al., 2023; Yismaw et al., 2025; Peng et al., 2024). We evaluate, post-training, the standard deviation of the error for each data dimension (i.e., cells or pixels), using different test data samples $i$ corrputed with known noise, as $\sigma_{{\hat{x}_{0}},i}^{(t)}\simeq\sqrt{\sum_{i=1}^{N}(\hat{x}_{0,i}^{(t)}-x_{0,i})^{2}/N}$. In fact, the error distribution depends on the spatial features of the input data, which is computable only through Hessian estimation (Efron, 2011; Boys et al., 2024; Hamidi and Yang, 2025) or using a network predicting conditional covariance (Peng et al., 2024, ; see Section 2.2). We assume error homoscedasticity, using the mean value of $\sigma_{{\hat{x}_{0}}}^{(t)}$ from all the pixels.
In agreement with Linde et al. (2017) and Friedli and Linde (2024) who studied approximations to intractable likelihoods arising in geophysics, we propagate the errors through the forward model using a first-order Taylor expansion $\mapsto\mathcal{F(\textbf{x})}$ x ) around $\hat{\textbf{x}}_{0}^{(t)}$,
@@
This is the main adaptation introduced by CDPS. We further observe another source of error in the original DPS implementation on DDPM (Chung et al., 2023). At each step, the authors use the likelihood score to obtain the conditional denoised image as $\textbf{x}_{t-1}=\textbf{x}_{t-1}^{\prime}-\rho\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})$, where $\textbf{x}_{t-1}^{\prime}$ is the denoised image considering the prior score, obtained with Eq. 8 as is. However, this is inconsistent with the appliaction of Bayes theorem (e.g., Eq. 10; Song et al., 2021b), according to which Eq. 8 in the conditioning case should be
$\textbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\ \hat{\epsilon}_{\theta}^{(t)}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\hat{\epsilon}_{\theta}^{(t)}+\sigma_{t}\epsilon_{t}.$ (20)
where $\hat{\epsilon}_{\theta}^{(t)}=\epsilon_{\theta}^{(t)}-\sqrt{1-\bar{\alpha}_{t}}\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})$, following the approach for guided diffusion modeling proposed by Dhariwal and Nichol (2021). In practice, even if one does not consider the error in the approximation $\hat{\textbf{x}}_{0}^{(t)}$, the likelihood score $\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})$ should be rescaled to the diffusion noise magnitudes rather than a fixed arbitrary $\rho$. We suggest that this error may be an additional cause of instability of the DPS method and provide an implementation of the CDPS for DDPM and DDIM in the Appendix C, implementing this correction.QUAQUA The CDPS algorithm for score-based diffusion is given in Algorithm 1, specifically within the EDM diffusion framework (Karras et al., 2022). Further details on EDM are provided in Section 2.3.
@@
1:Require: trained network $S_{\theta}(-)$, conditioning data d, forward operator $\mathcal{F}$, maximum number of time steps $N$, noise schedule $\sigma_{i}$, limits $\sigma_{max}$ and $\sigma_{min}$
@@ 2.3 Multivariate diffusion modeling
For posterior sampling (Eq. 10), we substitute the score with $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t}|\textbf{d})$ using our proposed likelihood approximation (Eq. 19). Sampling is performed evaluating the ODE at noise levels following the schedule (Karras et al., 2022)
@@
where $\sigma_{max}$ and $\sigma_{min}$ are the maximum and minimum noise levels in the corrupted data, $i$ is the step number, $i\in\{1,...,N\}$, and $\rho$ is an exponential factor. In agreement with the authors original choices, $\sigma_{max}=80$, $\sigma_{min}=0.002$ and $\rho=7$ provided the best performance; the motivation of these parameter choices is explained by Karras et al. (2022). The ODE is evaluated using the Heun’s $2^{\mathrm{nd}}$ order method (Ascher and Petzold, 1998) (lines 23-36 in Algorithm 1).
The network $S_{\theta}(\textbf{x}_{t},t)$ is trained as a denoiser, using an improved training algorithm with noise-specific weights integrated in Eq. 7 proposed by Karras et al. (2022). The network architecture relies on the encoder-decoder structure of a U-Net (Ronneberger et al., 2015), as implemented by Dhariwal and Nichol (2021) for general image generation. We use a multi-head self-attention mechanism (Vaswani et al., 2017) at each of its three layers to capture global dependencies. In our implementation, each subsurface property type is represented by a separate channel. Categorical variables such as geological facies assume a single numerical variable within a single dimension. Similarly to Miele et al. (2024) and Miele and Azevedo (2024), the multivariate prior distribution $p(\textbf{x}_{0})$ is represented by a training dataset where each sample is a joint realization of two or more co-located subsurface properties. The data can be represented as a single large image (randomly sampled during training) (see, e.g., Laloy et al., 2018) or as separate samples.
@@ 2.4 Performance assessment
Prior modeling is assessed by comparing the unconditional subsurface realizations against the TI and realizations generated with two benchmark models, a GAN and a VAE, proposed by Miele et al. (2024) for an analogous case study. The statistical distance between the target and modeled distributions (either prior or posterior) were measured using the root-mean-square-error (RMSE) and the Kullback-Leibler divergence ($KL$) (Kullback and Leibler, 1951). The structural similarity of spatial features in the realizations was assessed using the structural similarity index (SSIM) (Wang et al., 2004) defined as
@@
where u and v define sliding windows of size $M\times M$, on two different images, $\mu$ and $\sigma$ are their corresponding mean and standard deviations, and $C$ is a constant. Following the original implementation (Wang et al., 2004) and previous related work (e.g., Levy et al., 2023; Miele et al., 2024), we set $M=7$, $C_{1}=0.01$ and $C_{2}=0.03$. The SSIM ranges between -1 and 1, with 1 indicating a perfect match between images. Further metrics concerning the reproduction of facies volume fractions, two-points statistics (i.e., variograms), and lateral continuity (shape) of geological features, are used in the Appendix A.
@@
A WRMSE $\leq 1.1$ is chosen to indicate convergence in terms of data misfit as the residuals have a similar statistic as the observational noise. A WRMSE $<1$ may indicate overfitting the noisy data. Moreover, the predictive power of the inferred pdf is assessed using the logarithmic scoring rule (logS) (Good, 1952), computed as logS$(\hat{p},\textbf{d})=-\log\hat{p}(\textbf{d})$, where $\hat{p}$ is the inferred pdf and d is the real data. A lower value of logS corresponds to a better approximation. The proposed CDPS is compared to the original DPS (Chung et al., 2023) and to a sampling method based on Markov chain Monte Carlo (MCMC), to assess whether the inferred solutions are consistent. For MCMC, we use the preconditioned Crank–Nicolson algorithm (pCN) (Cotter et al., 2013), which is well-suited for high-dimensional problems. In this method, each new proposal $\textbf{x}_{new}$ is defined by $\textbf{x}_{new}=\sqrt{1-\beta^{2}}\textbf{x}_{old}+\beta\textbf{z}$, where $\textbf{x}_{old}$ represents the previous state, $\beta$ is a constant determining the step size, $\textbf{z}\sim\mathcal{N}(0,\textbf{C})$, and C is the prior covariance, determining the jump rate. The acceptance probability $\alpha$ takes the form $\alpha(\textbf{x}_{new},\textbf{x}_{old})=\min(1,(p(\textbf{d}|\textbf{x}_{new})/p(\textbf{d}|\textbf{x}_{old})))$. Convergence of the pCN samples is determined with the Gelman-Rubin diagnostic (Gelman and Rubin, 1992) using $\hat{R}\leq 1.2$ for all model parameters as a criterion. We further assess the exploration of the parameter space by computing within-chain autocorrelation of samples, that is measuring the covariance between two samples divided by the parameter’s variance.

### v2

@@ 2.2 Diffusion Posterior Sampling (DPS)
In Bayesian inversion, one is interested in characterizing or sampling from the posterior pdf, $p(\textbf{x}_{0}|\textbf{d})$, with the data vector d given by
@@
Unconditional modeling from $p(\textbf{x}_{0})$ can be achieved with a trained DM, using Eqs. 2 or 3. Following Bayes’ rule, sampling from the posterior pdf can in theory be achieved using the score function (Song et al., 2021b)
@@
where the first term can be approximated by the trained neural network $S_{\theta}(\textbf{x}_{t},t)$ and $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ is a noise-dependent likelihood score. In what follows, we will refer to $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$ as the prior score. If $S_{\theta}(\textbf{x}_{t},t)\simeq\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$ (Eq. 6), then $\hat{\textbf{x}}_{0}^{(t)}$ is derived using Tweedie’s formula (Eq. 5). Otherwise, it is directly obtained as $\hat{\textbf{x}}_{0}^{(t)}=S_{\theta}(\textbf{x}_{t},t)$ (Eq. 7). The likelihood score is given by an intractable integral
$p_{t}(\textbf{d}|\textbf{x}_{t})=\int{p(\textbf{d}|\textbf{x}_{0})p(\textbf{x}_{0}|\textbf{x}_{t})d\textbf{x}_{0}}.$ (11)
@@
In practice, the score in Eq. 12 is computed by backpropagating the log-likelihood approximations at each time step $t$. This approach is referred to as Diffusion Posterior Sampling (DPS), and was proposed by Chung et al. (2023) as a general inference framework for both linear and nonlinear inverse problems. Assuming uncorrelated and homoscedastic Gaussian data errors, Chung et al. (2023) use the likelihood
[…]
@@
[…]
This approximation does not account for errors in the $\hat{\textbf{x}}_{0}^{(t)}estimate$. In their implementation, Chung et al. (2023) propose a weight to the likelihood score, replacing $\frac{1}{\sigma_{\textbf{d}}^{2}}$ with a hyperparameter $\rho$, to account for the impact of errors in the likelihood approximation. In their accompanying code, the authors further replace the sum of square error in Eq. 14 with the root-mean-square error (RMSE), justifying this choice as a solution for increased stability (see open review in Chung et al., 2023). This arbitrary choice will drastically decrease the magnitude of the likelihood score and, hence, decrease the weight given to the data in the sampling process.
The conditional score approximation used within DPS results in a too low variance of the generated samples; it has been argued that the sampling process is rather associated to the maximization of a posterior than to posterior sampling (Xu et al., 2025). Improvements to this approximation should include the impact of errors in the denoiser ($\hat{x}_{0}^{(t)}$) (Song et al., 2023; Hamidi and Yang, 2025; Yismaw et al., 2025; Boys et al., 2024; Peng et al., 2024). Following Efron (2011), the denoiser’s variance can be computed using Tweedie’s formula for the second moment
@@
[…]
@@ 2.2.1 Corrected Diffusion Posterior Sampling (CDPS)
The proposed corrected DPS (CDPS), in which the likelihood formulation accounts for the denoising-step-dependent error $\epsilon_{\hat{\textbf{x}}_{0}}^{(t)}$ of the trained model. The $\epsilon_{\hat{\textbf{x}}_{0}}^{(t)}$ is the largest at early steps of the backward process and decreases with decreasing $\sigma_{\textbf{t}}$. The forward problem at time t can be written as (c.f., Eq. 9).
@@
Assuming an ideal denoiser, we assume $\epsilon_{\hat{\textbf{x}}_{0}}^{(t)}$ to be spatially uncorrelated and Gaussian-distributed (Song et al., 2023; Yismaw et al., 2025; Peng et al., 2024). After training, we evaluate the standard deviation of the error for each data dimension (i.e., cells or pixels), using different test data samples $i$ corrupted with known noise, as $\sigma_{{\hat{x}_{0}},i}^{(t)}\simeq\sqrt{\sum_{i=1}^{N}(\hat{x}_{0,i}^{(t)}-x_{0,i})^{2}/N}$. The error distribution depends on the spatial features of the input data, which is computable only through Hessian estimation (Efron, 2011; Boys et al., 2024; Hamidi and Yang, 2025) or using a network predicting conditional covariance (Peng et al., 2024, see Section 2.2). We assume error homoscedasticity, using the mean value of $\sigma_{{\hat{x}_{0}}}^{(t)}$ from all the pixels.
In agreement with Linde et al. (2017) and Friedli and Linde (2024) who studied approximations to intractable likelihoods arising in geophysics, we propagate the errors, $\sigma_{{\hat{x}_{0}},i}^{(t)}$, through the forward model using a first-order Taylor expansion $\textbf{x}\mapsto\mathcal{F(\textbf{x})}$ around $\hat{\textbf{x}}_{0}^{(t)}$,
@@
This is the main adaptation introduced by CDPS.
We report on another error in the original DPS implementation on DDPM (Chung et al., 2023). At each step, the authors use the likelihood score to obtain the conditional denoised image as $\textbf{x}_{t-1}=\textbf{x}_{t-1}^{\prime}-\rho\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})$, where $\textbf{x}_{t-1}^{\prime}$ is the denoised image considering the prior score, obtained with Eq. 8 as is. However, this is inconsistent with Bayes’ theorem (e.g., Eq. 10; Song et al., 2021b), according to which Eq. 8 in the conditioning case should be
$\textbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\ \hat{\epsilon}_{\theta}^{(t)}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\hat{\epsilon}_{\theta}^{(t)}+\sigma_{t}\epsilon_{t}.$ (20)
where $\hat{\epsilon}_{\theta}^{(t)}=\epsilon_{\theta}^{(t)}-\sqrt{1-\bar{\alpha}_{t}}\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})$, following the guided diffusion modeling approach proposed by Dhariwal and Nichol (2021). In practice, even if one does not consider the error in $\hat{\textbf{x}}_{0}^{(t)}$, the likelihood score $\nabla_{\textbf{x}_{t}}\log p(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})$ should be rescaled to the diffusion noise magnitudes rather than to a fixed arbitrary $\rho$. We suggest that this error may be an additional cause of instability of the DPS method; we provide an implementation of the CDPS for DDPM and DDIM in Appendix C, implementing this correction. The CDPS algorithm for score-based diffusion is given in Algorithm 1, specifically within the EDM diffusion framework (Karras et al., 2022). Further details on EDM are provided in Section 2.3.
@@
1:Require: trained network $S_{\theta}(-)$, conditioning data d, forward operator $\mathcal{F}$, maximum number of time steps NN, noise schedule $\sigma_{i}$, limits $\sigma_{max}$ and $\sigma_{min}$
@@ 2.3 Multivariate diffusion modeling
For posterior sampling (Eq. 10), we substitute the score with $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t}|\textbf{d})$ using our proposed likelihood approximation (Eq. 19). Sampling is performed by evaluating the ODE at noise levels given by the schedule (Karras et al., 2022)
@@
where $\sigma_{max}$ and $\sigma_{min}$ are the maximum and minimum noise levels in the corrupted data, $i$ is the step number, $i\in\{1,...,N\}$, and $\rho$ is an exponential factor. In agreement with the authors’ original choices, we found that $\sigma_{max}=80$, $\sigma_{min}=0.002$ and $\rho=7$ provided the best performance. The motivation behind these parameter choices is explained by Karras et al. (2022). The ODE is evaluated using the Heun’s $2^{\mathrm{nd}}$ order method (Ascher and Petzold, 1998) (lines 23-36 in Algorithm 1).
The network $S_{\theta}(\textbf{x}_{t},t)$ is trained as a denoiser, using an improved training algorithm with noise-specific weights integrated in Eq. 7 as proposed by Karras et al. (2022). The network architecture relies on the encoder-decoder structure of a U-Net (Ronneberger et al., 2015), as implemented by Dhariwal and Nichol (2021) for general image generation. We use a multi-head self-attention mechanism (Vaswani et al., 2017) at each of its three layers to capture global dependencies. In our implementation, each subsurface property type is represented by a separate channel. Categorical variables such as geological facies assume a single numerical variable within a single dimension. Similarly to Miele et al. (2024) and Miele and Azevedo (2024), the multivariate prior distribution $p(\textbf{x}_{0})$ is represented by a training dataset where each sample is a joint realization of two or more co-located subsurface properties. The data can be represented as a single large image (randomly sampled during training) (see, e.g., Laloy et al., 2018) or as separate samples.
@@ 2.4 Performance assessment
Prior modeling performance is assessed by comparing the unconditional subsurface realizations against the TI and realizations generated with two benchmark models, a GAN and a VAE, proposed by Miele et al. (2024) for an analogous case study. The statistical distance between the target and modeled distributions (either prior or posterior) were quantified using the root-mean-square-error (RMSE) and the Kullback-Leibler divergence ($KL$) (Kullback and Leibler, 1951). The similarity of spatial features in the realizations was assessed using the structural similarity index (SSIM) (Wang et al., 2004) defined as
@@
where u and v define sliding windows of size $M\times M$, on two different images, $\mu$ and $\sigma$ are their corresponding mean and standard deviations, and $C$ is a constant. Following the original implementation (Wang et al., 2004) and previous related works (e.g., Levy et al., 2023; Miele et al., 2024), we set $M=7$, $C_{1}=0.01$ and $C_{2}=0.03$. The SSIM ranges between -1 and 1, with 1 indicating a perfect match between images. Further metrics concerning the reproduction of facies volume fractions, two-points statistics (i.e., variograms), and lateral continuity (shape) of geological features, are presented in Appendix A.
@@
[…]